OpenStudy (anonymous):

WILL MEDALL & FANN The area of a rectangular painting is given by the trinomial x^2 - 5x - 36. What are the possible dimensions of the painting? Use factoring. a. (x+9) and (x-4) b. (x+9) and (x+4) c. (x-9) and (x-4) d. (x-9) and (x+4)

OpenStudy (anonymous):

@BoraSemiz

myininaya (myininaya):

It just wants you to factor the trinomial given

myininaya (myininaya):

What tow numbers multiply to be -36 and also add up to be -5

myininaya (myininaya):

two*

OpenStudy (anonymous):

-6 x 6?

OpenStudy (anonymous):

= -36 right?

myininaya (myininaya):

but I don't think -6+6 is -5

OpenStudy (anonymous):

what about -36 x 1 -5 x 1

myininaya (myininaya):

you know you are looking for 2 numbers that when you multiply give you -36 AND also when you those same numbers they give you -5 right?

myininaya (myininaya):

for example what two numbers multiply to be -54 and add up to be 3? well -6* 9=-54 and -6+9=3 so the two numbers I'm looking for in my example is -6 and 9

OpenStudy (anonymous):

oh okay..

OpenStudy (anonymous):

idk two numbers that equal -36 and -5....

myininaya (myininaya):

no no...

myininaya (myininaya):

you are looking for two numbers that multiply to be -36 and add up to be -5?

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhh okay..

myininaya (myininaya):

think about two numbers that multiply to be -36? -1(36)=-36 1(-36)=-36 2(-18)=-36 -2(18)=-36 3(-12)=-36 -3(12)=-36 4(-9)=-36 -9(4)=-36 so you couldn't tell me out of these possibilities which of the pair of factors can you add that give you -5?

myininaya (myininaya):

and we already discussed -6(6) -6(6)=-36 but -6+6 isn't -5

OpenStudy (anonymous):

4+-9

myininaya (myininaya):

yes

myininaya (myininaya):

so going back to my example let me extend my example: factor \[x^2+3x-54 \\ \text{ we found two factors of -54 that have product -54 and sum 3 } \\ \text{ those numbers were } -6 \text{ and } 9 \\ \text{ the factored form of } x^2+3x-54 \text{ is } (x-6)(x+9)\]

OpenStudy (anonymous):

so the factored form for x^2-5x-36 is (x+4) and (x-9)

myininaya (myininaya):

well the factors of x^2-5x-36 is (x+4) and (x-9) the factored form of x^2-5x-36 is (x+4)(x-9)

OpenStudy (anonymous):

oh okay i get it!!... thxx so much for ur help!! :)

myininaya (myininaya):

np if you get stuck at factoring things in the form x^2+bx+c you might want to try listing out the possibilities of c (hopefully c won't be too large because you know that can get pretty lengthy) and what I mean is you want two find two numbers that multiply to be c and also add up to be b so those numbers are m and n you are looking for m and n such that c=m*n and b=m+n so then you can write x^2+bx+c as (x+m)(x+n) multiplying (x+m)(x+n) out we have x^2+(m+n)x+mn as you see it is true comparing it to x^2+bx+c we need mn to be c and m+n to be b

OpenStudy (anonymous):

okaay.. thx again.. this really helped!! :)